OpenStudy (anonymous):

3x+6y=-6 -5x-2y=-14 solve system of equation elimination

OpenStudy (anonymous):

i got -3 as x

OpenStudy (anonymous):

multiply first equation with 5, it becomes \[15x+30y=-30 ... (I)\] multiply second equation with 3, it becomes \[-15x-6y=-42 .... (II)\] add \[I + II\] we have \[24y=-72 \implies y=-3\] put value of y in any of the equation to get the value of x. it comes out to be 4.

OpenStudy (anonymous):

solution is (4, -3)

OpenStudy (anonymous):

thankss

OpenStudy (anonymous):

your welcome. did you get it?

OpenStudy (anonymous):

cant we times -5x-2y=-1 by 3 ??

OpenStudy (anonymous):

for cross check you can put x=4 and y=-3 in original equation .. it must satisfy.

OpenStudy (anonymous):

I guess equation is \[-5x-2y=-14\]